Data: 25 25 26 27 28 30 30 31 32 35 35 38 16) Write a...

Data:
25
25
26
27
28
30
30
31
32
35
35
38

16) Write a sentence describing the shape of the distribution (skewed or symmetric).

17) How close together are your mean and median? Does this support your thoughts of skewness or symmetry for #16?

Use the empircal rule to answer the following:

1) Assume the mean is 29 and the standard deviantion is 2. Draw the curve, identify the mean at the center, count out the standard deviations from the center.

2) What is the probability someone has a BMI between 27 and 33?

3) What BMI's indicate the lowest 16% and highest 16%?
Answer #1

16) We need to find mean, standard deviation, and median and then using the following formula to get skewness of the given distribution.

3 * (m - me)

Where m = mean, me = median, sd = standard deviation

Let's use excel for calculation.

The formulae used on the above excel-sheet are as follows:

THe value of skweness is 0.123195 which is slightly greater than 0.

So we conclude that the given distribution is symmetric.

17) From the above output mean - median = 30.1667 - 30 = 0.1667

Since the difference between mean and median is very small, which support the above conclusion as the given distribution is symmetric.

Use the empircal rule to answer the following:

1) Assume the mean is 29 and the standard deviantion is 2. Draw the curve, identify the mean at the center, count out the standard deviations from the center.

2) What is the probability someone has a BMI between 27 and 33?

3) What BMI's indicate the lowest 16% and highest 16%?

o.3h 0.3h o135 o 135 o.0235 o. 0235 23 2 25 26 2 29 2 30 31 32 33 3 35 empini cof ue =0.815 3) from graph P [x <271 =1-0,5-o.35=o.16 the therefore BMIs indicetes the lowest é ond 8MZs indicates the highest 16. is 3